OpenStudy (anonymous):

Find the domain of the function f(x)=x-5/2x-1

OpenStudy (zzr0ck3r):

\(f(x)=\frac{x-5}{2x-1}\) ?

OpenStudy (anonymous):

the domain is all real numbers

OpenStudy (zzr0ck3r):

no its not

OpenStudy (anonymous):

yes it is

OpenStudy (zzr0ck3r):

and dont just give answers

OpenStudy (zzr0ck3r):

no its not, please stop....

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

@heatherjoy The only problem with rational functions composed of polynomials is that the denominator CANT equal to \(0\). SO in this situation \(2x-1\ne0\) what does this tell us that x CANT be?

OpenStudy (anonymous):

just giving me the answer doesn't help. i don't know how you got it and i have to show my work

OpenStudy (zzr0ck3r):

do you understand what I wrote?

OpenStudy (anonymous):

nope

OpenStudy (zzr0ck3r):

so you have \(\frac{x-5}{2x-1}\) correct

OpenStudy (zzr0ck3r):

?

OpenStudy (anonymous):

i can't see the top part of your post

OpenStudy (zzr0ck3r):

im confused..

OpenStudy (zzr0ck3r):

there is no top part, I just put space there

OpenStudy (anonymous):

is that a 5? it looks like a b

OpenStudy (zzr0ck3r):

\(\huge \frac{x-5}{2x-1}\)

OpenStudy (anonymous):

ok, yes

OpenStudy (zzr0ck3r):

so the only problem here is that we cant divide by \(0\)

OpenStudy (anonymous):

right

OpenStudy (zzr0ck3r):

in other words \(\huge 2x-1\ne 0\)

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

now treat that \(\ne \) like it was an \(=\) and solve for x \(2x-1\ne 0 \implies 2x\ne 1 \implies x\ne \frac{1}{2} \)

OpenStudy (zzr0ck3r):

so our domain is \(\{x\ |\ x\ne\frac{1}{2}\}\)

OpenStudy (zzr0ck3r):

or in words All \(x\) such that \(x\) is NOT \(\frac{1}{2}\)

OpenStudy (anonymous):

what is that fraction? i'm sorry. i can't see it

OpenStudy (zzr0ck3r):

\(\huge \frac{1}{2}\)

OpenStudy (zzr0ck3r):

\(\huge 2x-1\ne 0 \implies 2x\ne 1 \implies x\ne \frac{1}{2}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you tell me wht the steps are then?

OpenStudy (anonymous):

this is showing me what not to do

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i don't understand

OpenStudy (zzr0ck3r):

those are all the steps this is how I would answer this question: \(\huge 2x-1\ne 0 \implies 2x\ne 1 \implies x\ne \frac{1}{2}\) so our domain is \(D=\{x \ | \ x \ne \frac{1}{2}\}\)

OpenStudy (anonymous):

oh. wow. i'm really confused

OpenStudy (zzr0ck3r):

asking for the domain is the same thing as asking \(\huge\text{what is x allowed to be}?\)

OpenStudy (anonymous):

an integer

OpenStudy (zzr0ck3r):

nono

OpenStudy (zzr0ck3r):

read the whole thing

OpenStudy (zzr0ck3r):

when you ask for the domain of a function you are asking what is x allowed to be, then we notice that x can be anything we want as long as we dont divide by 0. The way you say that "mathematically" is \(\huge 2x-1\ne 0\)

OpenStudy (anonymous):

ive never seen this before. it looks like it's telling me x isn't allowed to be those things. i've never seen that symbol

OpenStudy (zzr0ck3r):

ok you dont have to write it like that, you can just say all x except 1/2

OpenStudy (anonymous):

i'm new to algebra, sorry

OpenStudy (zzr0ck3r):

its no problem

OpenStudy (zzr0ck3r):

most of the time we have to list domain in terms in what it is not allowed to be, because we cant list out every real number (there are infinite of them)

OpenStudy (anonymous):

ohhhhh

OpenStudy (zzr0ck3r):

so instead of saying what it can be, we say it can be anything except some certain numbers

OpenStudy (anonymous):

ok that makes sense

OpenStudy (zzr0ck3r):

so we decided that the denominator CANT be 0 our denominator is \(2x-1\)

OpenStudy (anonymous):

so the answer is what you wrote, simply because x isn't allowed to be anything else in the equation?

OpenStudy (zzr0ck3r):

our denominator CANT be 0 \(\downarrow\) \(\huge 2x-1\ne 0\)

OpenStudy (anonymous):

ok, so this is the beginning of solving?

OpenStudy (zzr0ck3r):

if we solve this we get that \(x\) cant be 1/2

OpenStudy (anonymous):

truthfully, i don't knw what a domain is, or what the problem is even asking me to do

OpenStudy (zzr0ck3r):

1/2 is the ONLY thing that will make the denominator 0. thus our domain is all real numbers EXCEPT 1/2

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

domain is just set of numbers we are allowed to plug into the function and get an answer that makes sense the domain of \(f(x)=\frac{1}{x}\) is all numbers EXCEPT 0 the domain of \(f(x) = 3x + 4\) is all real numbers, because there is no chance to divide by 0 the domain of \(f(x) = \frac{x+3}{x-7}\) is all real number EXCEPT 7, because x=7 makes the denominator be 0

OpenStudy (anonymous):

ok, so my answer is just that then? in this instance

OpenStudy (anonymous):

what about the 2?

OpenStudy (zzr0ck3r):

so we are only concerned with when 2x-1 is equal to 0 so we see when that happens \(\huge 2x-1=0 \implies 2x=1\implies x=\frac{1}{2}\)

OpenStudy (anonymous):

so is this a first step? or the only step in this problem?

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

are you still there?

OpenStudy (zzr0ck3r):

look up to where I said this is how I would answer the problem.